A bat strikes a 0.160 kg baseball. Just before impact, the ball is travelling horizontally to the right at 45.0 m/s, and it leaves the bat traveling to the left at an angle of 36.9 above the horizontal with at speed of 60.0 m/s. If the ball and the bat are in contact for 1.75 ms, (a) find the components of the


impulse of the net force on the ball. (b) What is the magnitude and direction of the average force on the ball?

1
Expert's answer
2022-04-21T13:02:36-0400

"v=v_1+v_2\cos\alpha=93~\frac ms,"

a)

"p=mv=15~\frac{kg\cdot m}s,"

b)

"F=\frac pt=8.5~kN."